OpenStudy (anonymous):

Factor Completely. c^3+22c^2+121c

OpenStudy (anonymous):

factor out the gcf first then see if the resulting trinomial is also factorable.

OpenStudy (anonymous):

c(c+11)^2

OpenStudy (anonymous):

\[c(c^2+22c+121)=c(c^2+22c+121)=c(c+11)^2\]